OpenStudy (anonymous):

what is the solution to the proportion? 3y-8/12= y/5

OpenStudy (amoodarya):

put variables in left side and numbers on the other side simplify then by divide to coefficient of variable you have the answer

OpenStudy (amoodarya):

\[3y-\frac{ 8 }{ 12 }=\frac{ y }{ 5 }\] is it ?

OpenStudy (amoodarya):

\[3y-\frac{ y }{ 5 }=\frac{ 8 }{ 12 }\] simplify

OpenStudy (triciaal):

(3y-8)/12 = y/5 lcd of 12 and 5 = 60 multiply each side by 60 5(3y-8) = 12y distribute , group like terms rewrite to get y=